A student studying the weather for d days observed that
1) it rained on 7 days , morning OR afternoon
2) When it rained in afternoon it was clear in morning
3) there were 5 clear afternoons
4) there were 6 clear mornings.
Then d equals

Answer by richwmiller(17219)

d=9
d--1-2-3-4-5-6-7-8-9-10-11
am-c-r-c-r-c-r-c-c-c------
pm-r-c-r-c-r-c-r-c-c------

((5+6)+7)/2=18/2=9
add partial clear 5+6=11 and 7 rainy days for a total of 18 ms and pms
divide 18/2 to get 9 days